# Evolving Symbolic 3D Visual Grounder with Weakly Supervised Reflection

### v2

@@ C.1.1 In Context Example Selection
The graph used for select in context example for relation encoder generation is shown in Figure 6.
@@
The detailed execution algorithm is presented in Algorithm 2, utilizing the precomputed category features and relation features. The Execute function runs recursively to compute the $\texttt{logits}\in\mathbb{R}^{N}$ ($N$ is the number of objects).
@@ Appendix D More Quantitative Results
We show evaluation results in NS3D(Hsu et al., 2023) in Table 4 . NS3D can only learn concepts (e.g. relation name, category name) from training set and its parsing results of Nr3D contains more than 5,000 concepts, resulting a long-tailed problem. So it selects a subset containing 1,041 test examples, which only contains the same concepts as Sr3D, the dataset it is mainly trained. On NS3D subset, EaSe have a large improvement than Ns3D (more than 7%), which shows the advantage of EaSe for processing natural grounding tasks.
@@
In Table 5, we show the evaluation results on NS3D, a subset of Nr3D, proposed by Hsu et al. (2023). Even not using training data of Sr3D, EaSe still achieves comparable performance with Hsu et al. (2023) on both setting (w/ and w/o GT labels). We show evaluation result on Sr3D, a subset of Achlioptas et al. (2020) in Table 5. In the standard setting, If using GT object labels, accuracy of our method (w/o VLM) on Sr3D is 95.3%, performance of NS3D and Fang et al. (2024) are 96.9% and 98.4%. So we believe that the bottleneck of Sr3D performance is object detection and classification rather than spatial relation understanding because its relation annotation is synthesized by relatively simple functions. So we mainly focus on natural benchmarks (Nr3D and Scanrefer) which have complex and real spatial relations.
@@ E.1 Effect of Code Optimization
We show the change between the initial response and the final code after multiple rounds of sampling and iterative refinement. By transitioning from a strict geometric overlap calculation to a continuous, exponential-based measure for both vertical and horizontal distances, the optimized code now captures nuances in “above” relationships more robustly. This improved formulation inherently handles scenarios where objects are close but not strictly overlapping, and it provides a more stable gradient for training. Consequently, the updated model passes all test cases by offering a smoother, more differentiable metric that better aligns with real-world spatial relations and passes more test cases.

### v3

@@ C.1.1 In Context Example Selection
The graph used for select in the context example for relation encoder generation is shown in Figure 6.
@@
[…]
@@ Appendix D More Quantitative Results
We show evaluation results in NS3D(Hsu et al., 2023) in Table 4 . NS3D can only learn concepts (e.g. relation name, category name) from the training set and its parsing results of Nr3D contain more than 5,000 concepts, resulting in a long-tailed problem. So it selects a subset containing 1,041 test examples, which only contains the same concepts as Sr3D, the dataset it is mainly trained on. On the NS3D subset, EaSe has a large improvement over Ns3D (more than 7%), which shows the advantage of EaSe for processing natural grounding tasks.
@@
In Table 5, we show the evaluation results on NS3D, a subset of Nr3D, proposed by Hsu et al. (2023). Even not using training data of Sr3D, EaSe still achieves comparable performance with Hsu et al. (2023) on both settings (w/ and w/o GT labels). We show evaluation results on Sr3D, a subset of Achlioptas et al. (2020) in Table 5. In the standard setting, if using GT object labels, the accuracy of our method (w/o VLM) on Sr3D is 95.3%, the performance of NS3D and Fang et al. (2024) are 96.9% and 98.4%. So we believe that the bottleneck of Sr3D performance is object detection and classification rather than spatial relation understanding because its relation annotation is synthesized by relatively simple functions. So we mainly focus on natural benchmarks (Nr3D and Scanrefer) which have complex and real spatial relations.
@@ E.1 Effect of Code Optimization
We show the change between the initial response and the final code after multiple rounds of sampling and iterative refinement. By transitioning from a strict geometric overlap calculation to a continuous, exponential-based measure for both vertical and horizontal distances, the optimized code now captures nuances in “above” relation more robustly. This improved formulation inherently handles scenarios where objects are close but not strictly overlapping, and it provides a more stable gradient for training. Consequently, the updated model passes all test cases by offering a smoother, more differentiable metric that better aligns with real-world spatial relations and passes more test cases.